OpenStudy (anonymous):

4x 2 y/2xy 3 simplify

OpenStudy (luffingsails):

\[\left( \frac{ 4x ^{2}y }{ 2xy^3 } \right)\] Like this?

OpenStudy (luffingsails):

@oaklandaslopez

OpenStudy (luffingsails):

If so, then the answer is: \[\frac{ 2x }{ y^2 }\]

OpenStudy (anonymous):

Your good. Thanks!

OpenStudy (luffingsails):

Sure.